# Criticality at work: scaling in the mouse cortex enhances performance

### v1

@@ Abstract
The critical brain hypothesis posits that neural systems operate near a phase transition, optimising the transmission, storage, and processing of information. While scale invariance and non-Gaussian dynamics — hallmarks of criticality — have been observed in brain activity, a direct link between criticality and behavioural performance remains unexplored. Here, we use a phenomenological renormalisation group (PRG) approach to examine neuronal activity in the primary visual cortex of mice performing a visual recognition task. We show that non-trivial scaling in neuronal activity is associated with enhanced task performance, with pronounced scaling observed during successful task completion. When rewards were removed or non-natural stimuli presented, these scaling signatures diminished. These results suggest that critical dynamics in the brain may be crucial for optimising behavioural outcomes, offering new insights into the functional role of criticality in cortical processing.
keywords:
brain criticality, renormalisation group, task performance, cortical dynamics
Introduction
@@
The idea that living systems self-organise in ways that promote efficiency, resilience, and adaptation has captivated scientists across multiple fields [1, 2, 3, 4, 5]. In neuroscience, this notion is encapsulated by the critical brain hypothesis [6, 7, 8, 9, 10, 11, 12, 13], which posits that the brain operates near a phase transition, a tipping point between two different states. This regime would be able to maximise information transmission, information storage and dynamic range in response to incoming stimuli [14, 15, 16, 7, 17, 18, 19, 20, 21], leading to debates about the possible clinical implications of the hypothesis [22, 23, 24, 25, 26].
@@
One of the hallmark features of the critical point where a phase transition occurs is scale invariance: the system’s activity has no characteristic length or time. Over two decades ago, this property was initially found in the self-similar temporal structure of human electroencephalogram (EEG) data [27] and in the scale-free nature of neuronal avalanches — bursts of electrical activity — in cortical slices [6]. Since then, a growing body of evidence supporting the existence of the putative phase transition has been gathered in different experimental setups [28, 29, 30, 31, 32, 33, 34, 35].
@@
Recently, a novel phenomenological renormalization group (PRG) method was introduced [36, 37], offering a fresh approach to studying scale-invariant behaviour in complex systems. Inspired by the renormalization group (RG) — the gold standard for understanding phase transitions in theoretical models [38, 39, 40, 41] — the PRG similarly reveals non-trivial scale invariance at non-Gaussian fixed points. It does so by tracking how observables evolve across scales as the system undergoes recursive coarse-graining. However, unlike the RG, the PRG is model-free and directly applicable to high-dimensional data from complex systems, making it particularly useful in neuroscience, where experimental advancements now enable the simultaneous recording of thousands of neurons [42]. The effectiveness of the PRG method has been validated through various studies, ranging from neural activity in C. Elegans to the whole human brain [37, 43, 44, 45, 46].
@@
Despite the increasing methodological sophistication in the methods to assess scale invariance in neuronal data, a glaring issue remains. Namely, that a link between criticality and task performance — the alleged reason why critical dynamics should be found in the brain in the first place — has never been found. Here we bridge this gap.
Results
Since our goal is to understand the relation between brain criticality and behaviour, we analysed neuron spiking activity from the primary visual cortex (V1) of mice while they performed a visual task [47] (Fig. 1a-c). This activity was subdivided into 30 s windows (Methods 1). For each window, we calculated the covariance matrix $C_{ij}$ comprising all pairs of recorded neurons. By diagonalising this matrix we obtained a set of eigenvectors from which projectors were created. When projecting the activity of each neuron over all eigenvectors, one recovers the original time series. By progressively selecting fewer components corresponding to the largest eigenvalues for projection, we generated an increasingly coarse-grained representation of the neural activity (a procedure akin to discarding faster modes in momentum space in standard RG) [36] (Fig. 1d). We tracked the normalised distribution of this activity as we chose fewer covariance eigenvectors. This procedure creates a flow of activity distributions (Fig. 1e-f) that pushes non-critical behaviour towards a trivial (Gaussian) fixed point. For each window, we used the kurtosis $\kappa$ of the distribution in the final coarse-graining step as a metric to quantify the deviation from the Gaussian distribution.
@@ Results
Figure 1: The phenomenological renormalisation group applied to mouse brain data. a, In the first part (Active) of the Allen Brain Observatory Visual Behavior Neuropixels experiment [47], natural images are shown in 250-ms flashes and the mouse needs to recognise a change of images in order to receive a reward. In sequence, the screen shows non-natural images such as Gabors, along with black and white flashes and a grey background. In the third part (Passive), the natural images are repeated in the same order as in the first part, but the mouse can no longer lick to be rewarded. b, Frontal slice of a mouse brain with the primary visual cortex (V1) highlighted in bright green (left) (credits to Allen Institute for Brain Science https://atlas.brain-map.org) and scheme of a Neuropixels 1.0 probe (right) used in the data collection. c, Raster plot of spiking activity for a 30-s window. d, Progressive coarse-graining of a unit time series, corresponding to projections onto a decreasing number of eigenvectors of the covariance matrix ($N,N/2,N/4,...$). The variance of the activity is normalized to one. e, Schematic representation of the renormalisation flow created by the coarse-graining procedure. Time windows that are further from criticality, such as those from spike-shuffled surrogate data, converge more rapidly to the Gaussian fixed point. f, In the final step of the coarse-graining procedure, windows closer to criticality (blue) retain their non-Gaussian shape, with higher kurtosis, while others (red) come closer to a Gaussian distribution (dashed line), with lower kurtosis.
@@
Certain time windows exhibit clear signatures of non-trivial dynamics, as expected at a critical point, with kurtosis values significantly higher than those in the spike-shuffled (surrogate) data. In contrast, other windows show results similar to surrogate data (Fig. 1f). Across 62 sessions (from 44 mice), we observed a broad range of kurtosis values, overlapping with the distribution from the surrogate data (Fig. 2a). As we depart from the Gaussian distribution, the distinction between windows with trivial and non-trivial statistics is arbitrary. We exploit this arbitrariness by establishing a kurtosis threshold $\kappa_{th}$: a window with $\kappa>\kappa_{th}$ is deemed non-trivial, with scaling properties; otherwise, it reflects Gaussian activity without scaling (Fig. 2a). We allow $\kappa_{th}$ to assume any value greater than zero.
@@
Figure 2: Improved task performance is linked to higher kurtoses. a, Distribution of kurtosis values at the final coarse-graining step (projections onto $N/16$ covariance eigenvectors) across time windows from all subjects (aggregated data). Unlike the surrogate data, the kurtosis of real data span a wider range of values. Windows with kurtosis values above (below) a tunable threshold ($\kappa_{th}$) are labelled as scaling (no-scaling) windows. b, The performance of real data is compared to data where hits and misses are shuffled (trial shuffled). c, The hit rate calculated across scaling windows is higher than both no-scaling windows and trial-shuffled data for all kurtoses exceeding the maximum value obtained for surrogate data $\kappa_{\text{max}}^{\text{surr}}$ (shaded area). Dashed line shows the kurtosis distribution. d, The relative performance gain is calculated from the difference between scaling and no-scaling curves, showing up to $24\%\pm 3\%$ increase in performance.
@@
The experiment has three phases. During the first phase, the mice perform a task of image change recognition every few seconds and are rewarded for correct responses (hits, Fig. 2b[47]. For any value of $\kappa_{th}$ above surrogate values, the average hit rate was consistently higher when calculated over scaling windows (Methods 1.1) than over no-scaling windows (Fig. 2c). We have verified that higher kurtosis was associated to better performance (Fig. 2c), with relative performance gain reaching up to $24\%\pm 3\%$ (Fig. 2d). As a control, we confirmed that the results were better than those obtained when hits and misses were shuffled within each session (trial-shuffled data, Fig. 2b). Time windows of 10 s and 60 s were also analysed and showed similar results, associating higher kurtosis with better performance (SI Fig.4).
@@
Figure 3: Comparison of the three phases of the experiment. a, Fraction of windows above a given kurtosis threshold ($\kappa_{th}$) in the aggregated data (complementary cumulative distribution function). The active phase remains consistently above the other phases for any choice of $\kappa_{th}$, indicating a less Gaussian-like behaviour. The shaded area corresponds to kurtoses of surrogate data. (b-e), comparison of scaling exponents ($\alpha$, $\beta$, $\mu$ and $z$) during the three phases. Bars (lines) represent the mean values (standard deviations) across sessions. Lighter shades depict results from surrogate data. Dashed lines correspond to the trivial values of exponents $\alpha$, $\beta$ and $z$. Triple asterisks (*) indicate a $p$-value $<0.001$ for the statistical comparison between two distributions utilising the Mann-Whitney U test.
@@
In the passive phase, the same images are shown, but there is no task to be performed or reward to be gained (the lick port is removed). During this phase, the distribution of kurtoses across windows got closer to surrogate values than in the active phase. For any $\kappa_{th}$, scaling windows were more abundant in the active phase than in the passive phase (Fig. 3a). Before the passive phase, mice were presented with non-natural images, such as Gabor patches, black and white screen flashes, or a grey background [47]. In this control condition, the shift towards surrogate levels in the kurtosis distribution was even more pronounced, and scaling windows became less likely to occur (Fig. 3a). These findings indicate that the process of performing a task drives an adaptation in the dynamics of the visual cortex, directing the system toward a non-Gaussian fixed point.
@@
We also calculated phenomenological renormalisation scaling exponents using a real-space approach to the coarse-graining method [37, 48] (Methods 1.3.2) (Fig. 3b-e). This process begins by calculating pairwise correlations from neuronal activity and progressively summing the time series of the most correlated pairs of neurons until there are none left unpaired. By iterating this process $k$ times, we obtain coarse-grained variables (or “clusters”) that are sums of $2^{k}$ spiking time series.
@@
Near a non-Gaussian fixed point, several quantities will scale with cluster size following non-trivial power laws, namely: the mean variance (exponent $\alpha$), minus the log of the probability of silence within a cluster ($\beta$), and the mean autocorrelation characteristic time ($z$). Additionally, the covariance eigenvalues scale with their rank with exponent $\mu$ [37, 48, 43, 44, 45] (Methods 1.3.2). For the active phase, we obtained $\alpha_{1}=1.41\pm 0.05$, $\beta_{1}=0.77\pm 0.04$, $\mu_{1}=0.73\pm 0.08$ and $z_{1}=0.3\pm 0.1$ (Fig. 3b-e). These scaling exponents show minimal change when calculated for kurtoses above $\kappa_{th}$, with only a weak dependence on this threshold (SI 6). During the passive phase we found $\alpha_{2}={1.37\pm 0.04}$, $\beta_{2}=0.78\pm 0.04$, $\mu_{2}=0.61\pm 0.05$, $z_{2}=0.36\pm 0.06$, with statistically significant differences with respect to the active phase for $\alpha$, $z$ and $\mu$ (Fig. 3b-e). These values align with those reported for awake mice in a different dataset [43]. Furthermore, for non-natural stimuli the exponents were $\alpha_{3}=1.28\pm 0.03$, $\beta_{3}=0.82\pm 0.03$, $\mu_{3}=0.38\pm 0.04$ and $z_{3}=0.4\pm 0.1$, which are all different from the phases with natural images (Fig. 3b-e).
Discussion
@@
We observed that non-Gaussian, scaling dynamics in the mouse primary visual cortex are closely linked with improved performance in a visual task, supporting the idea that critical dynamics underlie functional advantages in neural processing. These findings provide strong evidence for the critical brain hypothesis, which suggests that neural systems operate near a critical point to optimize information processing and may be essential for understanding brain function [2, 13].
@@
The kurtosis of the coarse-grained activity distribution offers a powerful means of assessing critical dynamics, with its flexibility — particularly the tunable threshold ($\kappa_{th}$) — enabling control over which windows are classified as ‘truly’ scaling. Our findings show that high-kurtosis windows consistently correspond to improved performance, marking these periods as indicative of a near-critical state. By increasing $\kappa_{th}$ to focus on only the highest kurtosis values, performance gains in scaling windows intensify, reaching up to $24\%\pm 3\%$ over non-scaling periods (Fig. 2c-d).
@@
The four scaling exponents also reflect the differences among the phases of the experiment. When animals are shown natural images and are actively performing the task, $\alpha$ and $\beta$ (“static exponents” [46]) deviate significantly from trivial values $\alpha=1$ and $\beta=1$ (Fig. 3b-c). In contrast, during a passive replay of natural images, the static exponents shift closer to trivial values, approaching this baseline even more closely when non-natural images are used as stimuli without any task involved. The “dynamic exponent” $z$ exhibited an inverse trend, diverging progressively from the trivial value $z=0$ across phases: increasing from active to passive and peaking in the non-natural image phase. This pattern is reminiscent of findings from a recently developed temporal renormalization group technique, which showed a deviation from critical behaviour during action compared to deep sleep [49]. Together, these findings provide new insights into the relationship between behaviour and criticality in the primary visual cortex, suggesting that task engagement and stimulus type modulate cortical scaling properties.
@@
Since pioneering experiments first reported neuronal avalanches in neural systems, numerous theoretical studies have proposed that the brain’s operation near criticality confers functional advantages, enhancing processing efficiency and adaptability. Our research bridges the gap between the critical brain hypothesis and its main flagship: that criticality underlies the brain’s efficient dynamics, facilitating optimized behaviour and cognition. Furthermore, our finding that scaling properties depend on the animal’s context opens new avenues for investigating how criticality might contribute to diverse brain functions and exploring the underlying mechanisms that sustain this state across varying cognitive demands and environmental challenges.
1 Methods
1.1 Data

### v2

@@ Abstract
The critical brain hypothesis posits that neural systems operate near a phase transition, optimizing the processing of information. While scale invariance and non-Gaussian dynamics—hallmarks of criticality—have been observed in brain activity, a direct link between criticality and behavioral performance remains unexplored. Here, we use a phenomenological renormalization group approach to examine neuronal activity in the primary visual cortex of mice performing a visual recognition task. We show that nontrivial scaling in neuronal activity is associated with enhanced task performance, with pronounced scaling observed during successful task completion. When rewards were removed or non-natural stimuli presented, scaling signatures diminished. These results suggest that critical dynamics in the brain is crucial for optimizing behavioral outcomes, offering new insights into the functional role of criticality in cortical processing.
I Introduction
@@
The idea that living systems self-organize in ways that promote efficiency, resilience, and adaptation has captivated scientists across multiple fields [1, 2, 3, 4, 5]. In neuroscience, this notion is encapsulated by the critical brain hypothesis [6, 7, 8, 9, 10, 11, 12, 13], which posits that the brain operates near a phase transition, a tipping point between two different states. This regime would be able to maximize information transmission, information storage, learning and dynamic range in response to incoming stimuli [14, 15, 16, 7, 17, 18, 19, 20, 21, 22, 23, 24, 25], leading to debates about the possible clinical implications of the hypothesis [26, 27, 28, 29, 30].
@@
One of the hallmark features of the critical point where a phase transition occurs is scale invariance: the system’s activity has no characteristic length or time. Over two decades ago, this property was initially found in the self-similar temporal structure of human electroencephalogram (EEG) data [31] and in the scale-free nature of neuronal avalanches—bursts of electrical activity—in cortical slices [6]. Since then, a growing body of evidence supporting the existence of the putative phase transition has been gathered in different experimental setups [32, 33, 34, 35, 36, 37, 38, 39, 40].
@@
Recently, a novel phenomenological renormalization group (PRG) method was introduced [41, 42], offering a fresh approach to studying scale-invariant behavior in complex systems. Inspired by the renormalization group (RG)—the gold standard for understanding phase transitions in theoretical models [43, 44, 45, 46]—the PRG similarly reveals nontrivial scale invariance at non-Gaussian fixed points. It does so by tracking how observables evolve across scales as the system undergoes recursive coarse-graining. However, unlike the RG, the PRG is model-free and directly applicable to high-dimensional data from complex systems, making it particularly useful in neuroscience, where experimental advancements now enable the simultaneous recording of thousands of neurons [47]. The effectiveness of the PRG method has been validated through various studies, ranging from neural activity in C. Elegans to the whole human brain [42, 48, 49, 50, 51].
@@
Despite the increasing methodological sophistication in the methods to assess scale invariance in neuronal data, a glaring issue remains. A link between criticality and task performance—the alleged reason why critical dynamics should be found in the brain in the first place—has never been found. Here we bridge this gap, showing that nontrivial scaling in the primary visual cortex of mice is associated with enhanced performance in a visual task.
@@
Figure 1: The phenomenological renormalization group applied to mouse brain data. (A) In the first part (Active) of the Allen Brain Observatory Visual Behavior Neuropixels experiment [52], natural images are shown in 250-ms flashes and the mouse needs to recognize a change of images in order to receive a reward. In sequence, the screen shows non-natural images such as Gabors, along with black and white flashes and a gray background. In the third part (Passive), the natural images are repeated in the same order as in the first part, but the mouse can no longer lick to be rewarded. (B) Frontal slice of a mouse brain with the primary visual cortex (V1) highlighted in bright green (left) (credits to Allen Institute for Brain Science) and scheme of a Neuropixels 1.0 probe (right) used in the data collection. (C) Raster plot of spiking activity for a 30-s window. (D) Progressive coarse-graining of a unit time series, corresponding to projections onto a decreasing number of eigenvectors of the covariance matrix ($N,N/2,N/4,\ldots$). The variance of the activity is normalized to 1.
Figure 2: The phenomenological renormalization group applied to mouse brain data. (A) Schematic representation of the renormalization flow created by the coarse-graining procedure. Time windows that are further from criticality, such as those from spike-shuffled surrogate data, converge more rapidly to the Gaussian fixed point. (B) In the final step of the coarse-graining procedure, windows closer to criticality (blue) retain their non-Gaussian shape, with higher kurtosis, while others (red) come closer to a Gaussian distribution (dashed line), with lower kurtosis (see Fig. S6 for the evolution of the distribution shape with kurtosis).
II Results
II.1 Performance increases with kurtosis
Since our goal is to understand the relation between brain criticality and behavior, we analyzed neuron spiking activity from the primary visual cortex (V1) of mice while they performed a visual task [53] [Figs. 1A-1C]. This activity was subdivided into 30-s windows (see the Appendix). For each window, we calculated the covariance matrix $C_{ij}$ comprising all pairs of recorded neurons. By diagonalizing this matrix, we obtained a set of eigenvectors from which projectors were created. When projecting the activity of each neuron over all eigenvectors, one recovers the original time series. By progressively selecting fewer components corresponding to the largest eigenvalues for projection, we generated an increasingly coarse-grained representation of the neural activity (a procedure akin to discarding faster modes in momentum space in standard RG) [41] [Fig. 1D]. We tracked the normalized distribution of this activity as we chose fewer covariance eigenvectors. This procedure creates a flow of activity distributions [Figs. 2A and 2B] that pushes noncritical behavior towards a trivial (Gaussian) fixed point. For each window, we used the kurtosis $\kappa$ of the distribution in the final coarse-graining step as a metric to quantify the deviation from the Gaussian distribution.
@@ II.1 Performance increases with kurtosis
Certain time windows exhibit clear signatures of nontrivial dynamics, as expected at a critical point, with kurtosis values significantly higher than those in the spike-shuffled (surrogate) data. In contrast, other windows show results similar to surrogate data [Fig. 2B]. Across 62 sessions (from 44 mice), we observed a broad range of kurtosis values, overlapping with the distribution from the surrogate data [Fig. 3A]. As we depart from the Gaussian distribution, the distinction between windows with trivial and nontrivial statistics is arbitrary. We exploit this arbitrariness by establishing a kurtosis threshold $\kappa_{\textrm{th}}$: a window with $\kappa>\kappa_{\textrm{th}}$ is deemed nontrivial, with scaling properties; otherwise, it reflects Gaussian activity without scaling [Fig. 3A]. We allow $\kappa_{\textrm{th}}$ to assume any value greater than zero.
@@
Figure 3: Improved task performance is linked to higher kurtoses. (A) Distribution of kurtosis values at the final coarse-graining step (projections onto $N/16$ covariance eigenvectors) across time windows from all subjects (aggregated data). Unlike the surrogate data, the kurtosis of real data span a wider range of values. Windows with kurtosis values above (below) a tunable threshold ($\kappa_{\textrm{th}}$) are labeled as scaling (no-scaling) windows. (B) The performance of real data is compared to data where hits and misses are shuffled (trial shuffled). (C) The hit rate calculated across scaling windows is higher than both no-scaling windows and trial-shuffled data for all kurtoses exceeding the maximum value obtained for surrogate data $\kappa_{\textrm{max}}^{\textrm{surr}}$ (shaded area). The dashed line shows the kurtosis distribution. (D) The relative performance gain is calculated from the difference between scaling and no-scaling curves, showing up to a $24\%\pm 3\%$ increase in performance.
@@
The experiment has three phases. During the first phase, the mice perform a task of image change recognition every few seconds, and they are rewarded for correct responses [hits, Fig. 3B] [52]. For any value of $\kappa_{\textrm{th}}$ above surrogate values, the average hit rate was consistently higher when calculated over scaling windows than over no-scaling windows [Fig. 3C]. We have verified that higher kurtosis was associated to better performance [Fig. 3C], with relative performance gain reaching up to $24\%\pm 3\%$ [Fig. 3D]. As a control, we confirmed that the results were better than those obtained when hits and misses were shuffled within each session [trial-shuffled data, Fig. 3B]. Time windows of 10 and 60 s were also analyzed and showed similar results, associating higher kurtosis with better performance (see the Supplemental Material).
@@
Figure 4: Comparison of the three phases of the experiment. (A) Fraction of windows above a given kurtosis threshold ($\kappa_{\textrm{th}}$) in the aggregated data (complementary cumulative distribution function). The active phase remains consistently above the other phases for any choice of $\kappa_{\textrm{th}}$, indicating a less Gaussian-like behavior. The shaded area corresponds to kurtoses of surrogate data. (B-E) Comparison of scaling exponents ($\alpha$, $\beta$, $\mu$ and $z$) during the three phases. Bars (lines) represent the mean values (standard deviations) across sessions. Lighter shades depict results from surrogate data. Dashed lines correspond to the trivial values of exponents $\alpha$, $\beta$ and $z$. Triple asterisks (*) indicate a $p$-value <0.001<0.001 for the statistical comparison between two distributions utilizing the Mann-Whitney U test.
II.2 Signatures of criticality change with context
@@
In the passive phase, the same images are shown, but there is no task to be performed or reward to be gained (the lick port is removed). During this phase, the distribution of kurtoses across windows got closer to surrogate values than in the active phase. For any $\kappa_{\textrm{th}}$, scaling windows were more abundant in the active phase than in the passive phase [Fig. 4A]. Before the passive phase, mice were presented with non-natural images, such as Gabor patches, black and white screen flashes, or a gray background [52]. In this control condition, the shift towards lower levels in the kurtosis distribution was even more pronounced, and scaling windows became less likely to occur [though most windows still remained with higher kurtosis than those of surrogate data, Fig. 4A]. These findings indicate that the process of performing a task drives an adaptation in the dynamics of the visual cortex, directing the system toward a non-Gaussian fixed point. Two additional analyses are available in the Supplemental Material. The first investigates to which extent the population firing rate relates to kurtosis and performance across experimental conditions. The second focuses on the possible effects of fatigue: in the second half of the active phase, we observe lower kurtoses than in the first half, but the relation between performance and kurtosis is preserved.
@@
We also calculated phenomenological renormalization scaling exponents using a real-space approach to the coarse-graining method [42, 54]. This process begins by calculating pairwise correlations from neuronal activity and progressively summing the time series of the most correlated pairs of neurons until there are none left unpaired. By iterating this process $k$ times, we obtain coarse-grained variables (or “clusters”) that are sums of $2^{k}$ spiking time series.
@@
Near a non-Gaussian fixed point, several quantities will scale with cluster size following nontrivial power laws, namely the mean variance (exponent $\alpha$), minus the log of the probability of silence within a cluster ($\beta$), and the mean autocorrelation characteristic time ($z$). Additionally, the covariance eigenvalues scale with their rank with exponent $\mu$ [42, 54, 48, 49, 50] (see Appendix .4). For the active phase, we obtained $\alpha_{1}=1.41\pm 0.05$, $\beta_{1}=0.77\pm 0.04$, $\mu_{1}=0.73\pm 0.08$ and $z_{1}=0.3\pm 0.1$ [Figs. 4B-4E]. These scaling exponents show minimal change when calculated for kurtoses above $\kappa_{\textrm{th}}$, with only a weak dependence on this threshold (see the Supplemental Material). During the passive phase we found $\alpha_{2}={1.37\pm 0.04}$, $\beta_{2}=0.78\pm 0.04$, $\mu_{2}=0.61\pm 0.05$, $z_{2}=0.36\pm 0.06$, with statistically significant differences with respect to the active phase for $\alpha$, $z$ and $\mu$ [Figs. 4B-4E]. These values align with those reported for awake mice in a different dataset [48]. Furthermore, for non-natural stimuli the exponents were $\alpha_{3}=1.28\pm 0.03$, $\beta_{3}=0.82\pm 0.03$, $\mu_{3}=0.38\pm 0.04$ and $z_{3}=0.4\pm 0.1$, which are all different from the phases with natural images [Figs. 4B-4E].
III Discussion
@@
We observed that non-Gaussian scaling dynamics in the mouse primary visual cortex is closely linked with improved performance in a visual task, supporting the idea that critical dynamics underlies functional advantages in neural processing. These findings provide strong evidence for the critical brain hypothesis, which suggests that neural systems operate near a critical point to optimize information processing and may be essential for understanding brain function [2, 13].
@@
The kurtosis of the coarse-grained activity distribution offers a powerful means of assessing critical dynamics, with its flexibility—particularly the tunable threshold ($\kappa_{\textrm{th}}$)—enabling control over which windows are classified as “truly” scaling. In essence, kurtosis acts as a de facto measure of the “distance to triviality.” Our findings show that as we consider only windows of progressively higher kurtosis (closer to the critical state), the corresponding performance calculated across those windows also rises. By increasing $\kappa_{\textrm{th}}$ to focus on only the highest kurtosis values, performance gains in scaling windows intensify, reaching up to $24\%\pm 3\%$ over non-scaling periods [Figs. 3C and 3D].
@@
The four scaling exponents also reflect the differences among the phases of the experiment. When animals are shown natural images and are actively performing the task, $\alpha$ and $\beta$ (“static exponents” [51]) deviate significantly from trivial values $\alpha=1$ and $\beta=1$ [Figs. 4B and 4C]. In contrast, during a passive replay of natural images, the static exponents shift closer to trivial values, approaching this baseline even more closely when non-natural images are used as stimuli without any task involved. The “dynamic exponent” $z$ exhibited an inverse trend, diverging progressively from the trivial value $z=0$ across phases: increasing from active to passive and peaking in the non-natural image phase. This pattern is reminiscent of findings from a recently developed temporal renormalization group technique, which showed a deviation from critical behavior during action compared to deep sleep [55]. Furthermore, the measured values of the exponent $\mu$ reinforce the idea that these differences are not purely input-driven [56]. Instead, the observed values align more closely with background activity (orthogonal to the input), as reported in a recent study [48]. Together, these findings provide new insights into the relationship between behavior and criticality in the primary visual cortex, suggesting that task engagement and stimulus type modulate cortical scaling properties.
The spread in both exponent and kurtosis values observed in Fig. 4 is consistent with a recent computational study using this technique [57]. That study reports that PRG produces a continuum of values for these quantities near a phase transition, with a distinct peak appearing only at the critical point. Taken together with our results, this suggests that the brain may operate not exactly at criticality, but rather in a slightly off-critical regime (such as a subcritical state), a view consistent with previous discussions in the literature [58].
@@
Since pioneering experiments first reported neuronal avalanches in neural systems, numerous theoretical studies have proposed that the brain’s operation near criticality confers functional advantages, enhancing processing efficiency and adaptability. Although earlier studies examined these effects in cortical slices [6], anesthetized rats [21], or explored the functional role of criticality in health-related contexts [59], no prior work has directly associated neuronal criticality with functional performance in a living, awake, and behaving animal, nor shown how these signatures of criticality change with task engagement. Though we are not able to provide a biophysical mechanism by which brain criticality enhances performance, our research establishes a phenomenological link between the critical brain hypothesis and its main flagship: that criticality underlies the brain’s efficient dynamics, facilitating optimized behavior and cognition. Furthermore, our finding that scaling properties depend on the animal’s context opens new avenues for investigating how criticality might contribute to diverse brain functions and exploring the underlying mechanisms that sustain this state across varying cognitive demands and environmental challenges.
Acknowledgements.
The authors acknowledge support from Brazilian agencies CNPq (Grants No. 140660/2022-4, No. 308703/2022-7, No. 314094/2023-7, No. 444500/2024-3 and No. 408389/2024-9), FACEPE (Grant No. APQ-1187-1.05/22), FINEP (Grant No. NeuroAssist 01.24.0124.00) and FADE/UFPE (Grant No. 64/2024). M.C. and D.M.C. designed the study. D.M.C. provided initial code for the analysis. G.G.C. refined the code, ran the analysis and made the plots. All authors discussed the results and wrote the paper.
DATA AVAILABILITY
The data that support the findings of this article are openly available [60].
Appendix: Methods and Materials
.1 Data